OpenStudy (anonymous):

If y varies inversely as x, and y = 10 when x = 7, find y for the x-value of 10.

OpenStudy (anonymous):

thank you :)

OpenStudy (aripotta):

first find k y = k/x 10 = k/7 k = 70

OpenStudy (aripotta):

then you can find y y = k/x y = 70/10 y = 7